If the distance covered by a body in nth second is given (4 + 6n)Then find the accelaration and initial speed If the distance covered by a body in nth second is given (4 +6n)Then find the accelaration and initial speed

Grade:9

1 Answers

Neel Patel
7 years ago
Snth=u+a/2(2n-1)
      =u-a/2+an.......(1)
 
Snth=(4+6n).....(2)
 
comparing both (1) and (2)
  an=6n
 a=6m/sec^2
 
 u-a/2=4
and
  u=7m/sec
 
hello! my name is neel.
 u like my ans?